The magnitude of L, for a d-electron, in one electron atomic system is?
1
Expert's answer
2020-06-03T12:01:53-0400

The orbital quantum number of d electron is "l=2"

The magnitude of total orbital quantum number for a d-electron in one electron atomic system is

"L=\sqrt{l(l+1)} \frac{h}{2 \pi}"

"=\sqrt{2(2+1)} \frac{h}{2 \pi}=\sqrt{6} \frac{h}{2 \pi}"